# Observational Auditing of Label Privacy

### v1

@@ 5 Observational attribute inference
5: Run attack $A$ with input $o_{1}=M(D^{0})$, $o_{2}=D^{b}$, and access to $\mathcal{D}$, $\mathcal{D}^{\prime}$.
@@
Similar to prior auditing papers (Mahloujifar et al., 2025b; Steinke et al., 2024) the adversary can choose to abstain from guessing on samples where it is least confident, to boost its positive likelihood ratio. The observational game can use the entire dataset as canaries (i.e., $m=n$).
@@
Obtaining approximate distributions. A key aspect in implementing the observational attribute inference attack is to produce the proxy distribution $\mathcal{D}^{\prime}$ from which the counterfactual partial records $y_{i}^{1}$ are sampled. One option is to train an additional model $M^{\prime}$ to predict the missing attribute(s). For label inference attacks, which are a special case of Algorithm 2, we sample counterfactual label $y_{i}^{i}$ from $\mathsf{Multinoulli}(M^{\prime}(x_{i}))$. In an online machine learning system, where the model trains incrementally as more recent data becomes available, one can use a prior model checkpoint as the model $M^{\prime}$ (and run the attack on the newer data). This eliminates the need for training any additional models, making our proposed label inference attacks very lightweight in terms of computational overhead and implementation complexity.
Auditing guarantees when $\mathcal{D}=\mathcal{D}^{\prime}$. We first establish auditing guarantees for the simpler case where $\mathcal{D}=\mathcal{D}^{\prime}$, i.e., when we can sample counterfactual records from the ground truth distribution. Our bounds are stated for auditing $f$-DP, following an analogous argument in Mahloujifar et al. (2025b), which gives the tightest DP auditing analysis and can be translated to the language of $({\epsilon},\delta)$. The auditing applies an evaluation procedure, Algorithm 3, described and analyzed in Appendix 9.
Theorem 5 (Auditing $f$-DP with no distribution shift).
@@
Let $M\colon(\mathcal{X}\times\mathcal{Y})^{*}\to\mathcal{Z}$ be a mechanism, $\mathcal{D}$ the data distribution, and $\mathrm{Sim}_{M,\mathcal{D}}$ the imputation-based simulator (Definition 2). Let $C=\sum_{i\in[m]}\mathbf{1}[b_{i}^{\prime}=b_{i}]$ be the total number of correct answers from the one-run observational attribute inference attack with no distribution shift (Algorithm 2, when $\mathcal{D}^{\prime}=\mathcal{D}$) for an adversary that makes $c^{\prime}$ guesses. If $M$ is f-SIM-DP and Algorithm 3 returns True on $(c^{\prime},c,M,f,\gamma=0.05)$, then $\Pr[C\geq c]\leq\gamma.$
Auditing guarantees under distribution shift. Now we consider the more general case when $\mathcal{D}\neq\mathcal{D}^{\prime}$. We state an information-theoretic bound, which depends on the total variation (TV) distance between $\mathcal{D}$ and $\mathcal{D}^{\prime}$. Intuitively, the larger the distance between the two distributions, the weaker the lower bound we can obtain on the privacy guarantee. The value $\tau$ in Theorem 6 is an upper bound on the adversary’s a priori (i.e., before having access to the target model) success probability in distinguishing whether a sample $(x,y^{b})$ is from $\mathcal{D}$ or $\mathcal{D}^{\prime}$.
Theorem 6 can also be stated in terms of the adversary’s ability to distinguish between $\mathcal{D}$ and $\mathcal{D}^{\prime}$ given its resource constraints. Such a bound is particularly meaningful for Label DP auditing, where we can assume that the adversary cannot distinguish $y\mid x$ from $\mathsf{Multinoulli}(M^{\prime}(x))$ if $M^{\prime}$ is the best classifier on $\mathcal{D}$ available to the adversary.
We state here a bound akin to Steinke et al. (2024), that is easier to interpret in the case of distribution shift. We present a bound on auditing $f$-DP under distribution shift in Appendix 9.
Theorem 6 (Auditing $({\epsilon},\delta)$-DP under distribution shift).
Let $M\colon(\mathcal{X}\times\mathcal{Y})^{*}\to\mathcal{Z}$ be a mechanism, $\mathcal{D}$ the data distribution, $\mathcal{D}^{\prime}$ the approximate distribution, and $\mathrm{Sim}_{M,\mathcal{D}^{\prime}}$ the imputation-based simulator. Let $C=\sum_{i\in[m]}\mathbf{1}[b_{i}^{\prime}=b_{i}]$ be the total number of correct answers from the one-run observational attribute inference attack (Algorithm 2). If $\mathcal{M}$ is $({\epsilon},\delta)$-$\mathrm{Sim}_{M,\mathcal{D}^{\prime}}$-DP and $\mathrm{TV}(\mathcal{D}|x_{i},\mathcal{D}^{\prime}|x_{i})\leq\tau$ for all $x_{i}\in D$, then
$\Pr_{S,D}[C\geq c\mid A(o_{1},o_{2})=b^{\prime}]\leq\Pr_{\hat{C}_{i}\sim\mathsf{Bernoulli}(\beta)}\Big[\sum_{i\in[m]}\hat{C}_{i}\geq c\Big]+n\delta$ (3)
where $\beta=\frac{e^{\epsilon}}{e^{\epsilon}+\frac{1-\tau}{1+\tau}}$.
A full theorem statement and its proof are deferred to Appendix 10.
@@ 6 Experiments on auditing Label DP
We validate our theoretical framework with experiments on two real representative datasets: CIFAR-10 (Krizhevsky, 2009) for image classification and Criteo (Criteo AI Lab, 2015) for tabular data with sensitive labels. For each dataset, we train classifiers using standard Label DP mechanisms and empirically evaluate the privacy guarantees using our audit procedure. We also evaluate our Label DP auditing technique for Randomized Response (Warner, 1965) on synthetic data. Our attack is implemented in the PrivacyGuard library (Melis et al., 2025). [^1] https://github.com/facebookresearch/PrivacyGuard
@@ 6.1 Label DP learning algorithms
The earliest approach to achieving Label DP is the Randomized Response (RR) mechanism (Warner, 1965). In RR, each training label is randomly replaced according to a fixed probability distribution before being shared with the learning algorithm. This randomization helps protect the privacy of individual labels. We briefly review several recent Label DP mechanisms that improve on RR.
Label Private One-Stage Training (LP-1ST, Ghazi et al. (2021)) Instead of using a fixed distribution as in RR, LP-1ST samples each training label $y_{i}$ from a learned prior distribution $P(y|X_{i})$. The prior can be estimated by observing the top-$k$ predictions from a pretrained model (either in domain or out-of-domain), restricting RR to the most probable labels. Alternatively, the training can be split into multiple stages, where an earlier model provides the prior for the next (LP-MST).
Private Aggregation of Teacher Ensembles with FixMatch (PATE-FM, Malek et al. (2021)) PATE-FM combines the FixMatch semi-supervised learning algorithm (Sohn et al., 2020) with private aggregation. Multiple teacher models are trained, each using all unlabeled examples and disjoint subsets of the labeled data. The predictions from these teachers are then aggregated in a differentially private manner using the PATE framework (Papernot et al., 2017) to train a student model.
Additive Laplace with Iterative Bayesian Inference (ALIBI, Malek et al. (2021)) ALIBI achieves Label DP by adding Laplace noise to the one-hot encoded labels (Ghosh et al., 2012), making the released labels differentially private. The model is trained on soft labels computed from noisy observations via Bayesian inference.
@@ 6.2 Attack implementation and adversarial strategy
For the Criteo dataset, where data is collected over 28 consecutive days, we train $M^{\prime}$ on Day 0 data and the target model on Day 1 data. While there may be some distributional shift between Day 0 and Day 1 data, we assume this is small, and that the adversary cannot distinguish between the true labels $y^{0}$ and the reconstructed labels $y^{1}$ (before having access to the model).
For CIFAR-10 experiments we randomly split the training data (n=50n=50K samples) into two. We train $M^{\prime}$ on the first half, and the target model $M$ on the second half.
We run the attack on m=200m=200K canaries for Criteo and m=10m=10K canaries for CIFAR-10. For the MIA experiments we use the same number of canaries for the non-members (taken from the test set).
@@
$s_{2}(x,y^{b})=\Pr[y^{0}\neq y^{1}]=1-M^{\prime}(x)[y^{b}].$
The final score combines the two components as $s(x,y^{b})=s_{1}(x,y^{b})\cdot s_{2}(x,y^{b})^{t}$ with a hyperparameter $t\geq 0$ that allows for weighting the two components separately. We use $t=2$, as $t>1$ gives tight lower bounds for RR. The adversary guesses on $c^{\prime}\%$ of samples with highest absolute scores. We sweep $c^{\prime}\in\{1,2,\dots,100\}$ and report the highest $\epsilon$} achieved at 95% confidence, averaged over 100 repetitions of the game (resamplings of counterfactual labels).
@@ 6.3 CIFAR-10 experiments
For CIFAR-10, we treat the image classes as sensitive labels. We train standard convolutional neural networks with varying privacy budgets ${\epsilon}\in\{1.0,10.0,\infty\}$ (see model accuracy in Table 4, Appendix 12). We then audit Label DP using our observational game and report results in Table 1.
@@
6.4 Criteo experiments
The Criteo dataset contains user click-through data with demographic information encoded as 13 numerical features and 26 categorical features. A binary label indicates whether the user clicked on the ad. The distribution of the labels is highly imbalanced, with only 3% of positives (clicks). The overall dataset contains over 4 billion click log data points over a period of 24 days. We followed the same setup in Wu et al. (2023) where 1 million data points are selected for each day. We divide the data into 80% for training, 4% for validation, and 16% for testing. Model performance is evaluated using the log-loss metric on the test set.
We train gradient boosting decision trees with the CatBoost library (Prokhorenkova et al., 2018) with varying privacy budgets ${\epsilon}\in\{0.1,1.0,2,0,4.0,8.0,\infty\}$ and evaluate label privacy using observational privacy auditing (Table 2). Table 5 (Appendix 12) shows model performance under different Label DP algorithms and values $\epsilon$}.
Table 2: Criteo. Auditing Label DP algorithms under different $\epsilon$} with $\delta=10^{-5}$. Similarly to Wu et al. (2023), for LP-1ST (domain prior) at ${\epsilon}\in\{0.1,1,2\}$ the training process did not produce meaningful outcomes.
@@ 6.5 Comparison with Existing Methods
We compare our observational auditing approach against traditional canary-based methods to demonstrate the effectiveness and practicality of our framework. More specifically, we evaluate against the lightweight difficulty calibration MIA in Watson et al. (2022), where membership scores are adjusted to the difficulty of correctly classifying the target sample. For each canary datapoint, we set the calibrated membership scores as the difference in the loss between the target model and the reference model $M^{\prime}$. Fig. 1 shows how our method achieves similar auditing results when compared to MIA on the CIFAR-10 and Criteo datasets.
@@
We leave for future work a comparison with more computationally intensive methods, such as Zarifzadeh et al. (2024), which require training multiple auxiliary (shadow) models to achieve state-of-the-art attack performances.
Figure 2: Auditing randomized response when the adversary guesses on 0.1% of samples. The counterfactual labels are generated either from the ground-truth distribution or a proxy distribution obtained from the predictions of logistic regression model.
@@ 6.6 Synthetic data and auditing randomized response
We empirically demonstrate the tightness of our Label DP auditing algorithm for Randomized Response using synthetic data. The distribution consists of $k$ balanced classes, each generated from a $5$-dimensional Gaussian with the same covariance but a shifted mean. We generate $n=10^{6}$ samples and experiment with $k\in\{2,5,10\}$. Counterfactual labels $y^{1}$ are generated using either the true distribution or the predictions of a logistic regression model. Figure 2 shows empirical epsilon lower bounds at $95\%$ confidence when the adversary makes $0.1\%$ non-abstaining guesses. We obtain tight lower bounds for ${\epsilon}\in[1,4]$. At lower epsilons the audit overestimates privacy loss due to a higher variance induced by a small number of guesses. Using more guesses at lower epsilon fixes the issue (Appendix 11). Obtaining tight lower bounds at very high epsilon is a limitation of current auditing methodology (Steinke et al., 2024; Mahloujifar et al., 2025b).

### v2

@@ 5 Observational attribute inference
5: Run attack AA with input $M(D^{0})$, $D^{b}$, and access to $\mathcal{D}$, $\mathcal{D}^{\prime}$.
@@
Similar to prior auditing papers (mahloujifar2025auditing; steinke2024privacy) the adversary can choose to abstain from guessing on samples where it is least confident, to boost its positive likelihood ratio. The observational game can use the entire dataset as canaries (i.e., $m=n$).
@@
Implementing the attack in practice, i.e., obtaining approximate distributions. A key aspect in implementing the observational attribute inference attack is to produce the proxy distribution $\mathcal{D}^{\prime}$ from which the counterfactual partial records $y_{i}^{1}$ are sampled. One option is to train an additional model $M^{\prime}$ to predict the missing attribute(s). For label inference attacks, which are a special case of Algorithm 2, we sample counterfactual label $y_{i}^{i}$ from $\mathsf{Multinoulli}(M^{\prime}(x_{i}))$. In an online machine learning system, where the model trains incrementally as more recent data becomes available, one can use a prior model checkpoint as the model $M^{\prime}$ (and run the attack on the newer data). This eliminates the need for training any additional models, making our proposed label inference attack very lightweight in terms of computational overhead and implementation complexity.
It is harder to operationalize Algorithm 2 for the case of MIA or attribute inference attacks, which require synthesizing entire samples or multiple attributes. In this case, diffusion models or generative adversarial models (GANs) can be used to generate the synthetic records, as in kazmi2024panoramia. However, the architecture of the generative model might differ from the architecture of the target model, thus requiring training an additional model. This presents a tradeoff in implementation complexity between training an additional model versus modifying training data pipelines to exclude data. In contrast, the observational label inference attack may use prior model checkpoints and requires no additional training.
Auditing guarantees under distribution shift. We now establish theoretical auditing guarantees that translate the attacker’s accuracy into a lower bound on the privacy parameters. Our bound depends on the total variation (TV) distance between $\mathcal{D}\mid x$ and $\mathcal{D}^{\prime}\mid x$. Intuitively, the larger the distance between the two distributions, the weaker the lower bound we can obtain on the privacy guarantee.
Theorem 5 (Auditing $f$-DP with distribution shift).
@@
Let $M\colon(\mathcal{X},\mathcal{Y})^{*}\to\mathcal{Z}$ be a mechanism, $\mathcal{D}$ the data distribution, $\mathcal{D}^{\prime}$ an approximate distribution, and $\mathrm{Sim}_{M,\mathcal{D}^{\prime}}$ the imputation-based simulator (Definition 2). Let $C=\sum_{i\in[m]}\mathbf{1}[b_{i}^{\prime}=b_{i}]$ be the total number of correct answers from the one-run observational attribute inference attack (Algorithm 2) for an adversary that makes $c^{\prime}$ guesses. Let $\mathrm{TV}(\mathcal{D}|x,\mathcal{D}^{\prime}|x)\leq\tau$ for all $x$ in the dataset $D$ and define $g\colon[0,1]\to[0,1]$ such that
$g(s)=f(\min(1,s+\tau)).$ (2)
If $M$ is $(\epsilon,\delta)$-SIM-DP with respect to $\mathrm{Sim}_{M,\mathcal{D}^{\prime}}$ and the auditing Algorithm 3 returns False on $(c^{\prime},c,M,g,\gamma)$, then $\Pr[C\geq c]\leq\gamma.$
Accounting for the distribution shift in practice. Theorem 5 requires that the bound $\tau$ be known to the challenger, which in practice can be hard to estimate. Not accounting for $\tau$ properly leads to an overestimation of the privacy leakage. Nonetheless, for the case of Label DP auditing we recommend using Theorem 5 with $\tau=0$ if the challenger is training the best available model for the prediction task. To elaborate, Theorem 5 can also be stated in terms of the adversary’s ability to distinguish between $\mathcal{D}$ and $\mathcal{D}^{\prime}$ given its knowledge and resource constraints. The value $\tau$ in Theorem 5 is an upper bound on the adversary’s a priori (i.e., before having access to the target model) success probability in distinguishing whether a sample $(x,y^{b})$ is from $\mathcal{D}$ or $\mathcal{D}^{\prime}$. In the case of Label DP auditing, if $M^{\prime}$ denotes the best classifier on $\mathcal{D}$ that the adversary can access, we may reasonably assume that the adversary is unable to distinguish the true conditional label distribution $y\mid x$ from the distribution $\mathsf{Multinoulli}(M^{\prime}(x))$.
Fig. 4 (Appendix 10) presents empirical impact of $\tau$ on the measured privacy loss on synthetic data. Higher values of $\tau$ correspond to lower measured privacy losses, due to the additive $\tau$ term in Eq. 2. While the the additive dependence can significantly degrade the measured empirical privacy loss, the effect can be mitigated by assuming stronger bounds on the distribution divergence (e.g., using the Kullback-Leibler divergence instead of the TV distance).
@@ 6 Experiments on auditing Label DP
We validate our theoretical framework with experiments on two representative datasets: CIFAR-10 (cifar10) for image classification and Criteo (criteo) for tabular data with sensitive labels. For each dataset, we train classifiers using standard Label DP mechanisms and empirically evaluate the privacy guarantees using our audit procedure. Appendix 12 shows similar results for a third large text dataset. We also evaluate our Label DP auditing technique for Randomized Response (warner1965randomized) on synthetic data.
@@ 6.1 Label DP learning algorithms
The earliest approach to achieving Label DP is the Randomized Response (RR) mechanism (warner1965randomized). In RR, each training label is randomly replaced according to a fixed probability distribution before being shared with the learning algorithm. This randomization helps protect the privacy of individual labels. We briefly review several recent Label DP mechanisms that improve on RR.
Label Private One-Stage Training (LP-1ST, google-labeldp) Instead of using a fixed distribution as in RR, LP-1ST samples each training label $y_{i}$ from a learned prior distribution $P(y\mid X_{i})$. The prior can be estimated by observing the top-$k$ predictions from a pretrained model (either in domain or out-of-domain), restricting RR to the most probable labels. Alternatively, the training can be split into multiple stages, where an earlier model provides the prior for the next (LP-MST).
Private Aggregation of Teacher Ensembles with FixMatch (PATE-FM, meta-labeldp) PATE-FM combines the FixMatch semi-supervised learning algorithm (sohn2020fixmatch) with private aggregation. Multiple teacher models are trained, each using all unlabeled examples and disjoint subsets of the labeled data. The predictions from these teachers are then aggregated in a differentially private manner using the PATE framework (papernot2016semi) to train a student model.
Additive Laplace with Iterative Bayesian Inference (ALIBI, meta-labeldp) ALIBI achieves Label DP by adding Laplace noise to the one-hot encoded labels (ghosh2012universally), making the released labels differentially private. The model is trained on soft labels computed from noisy observations via Bayesian inference.
@@ 6.2 Attack implementation and adversarial strategy
For the Criteo dataset, where data is collected over 28 consecutive days, we train $M^{\prime}$ on Day 0 data and the target model on Day 1 data. While there may be some distribution shift between Day 0 and Day 1 data, we assume this is small, i.e., ($\tau$ in Theorem 5 is 0) and that the adversary cannot distinguish between the true labels $y^{0}$ and the reconstructed labels $y^{1}$ without access to the model. For CIFAR-10 experiments we randomly split the training data (n=50n=50K samples) into two. We train $M^{\prime}$ on the first half, and the target model $M$ on the second half.
We run the attack on m=200m=200K canaries for Criteo and m=10m=10K canaries for CIFAR-10. One of the benefits of our observational attack is that it can be run with as many canaries as the size of the training set, eliminating a prior tradeoff for interventional attacks, where using more canaries gives tighter confidence intervals but decreases the size of the available training set. To use the same number of canaries as for MIA experiments, the number of canaries is set to the size of the test sets.
@@
$s_{2}(x,y^{b})$ $\mathrel{\triangleq}\Pr[y^{0}\neq y^{1}]=1-M^{\prime}(x)[y^{b}].$
The final score combines the two components as $s(x,y^{b})\mathrel{\triangleq}s_{1}(x,y^{b})\cdot s_{2}(x,y^{b})^{t}$ with a hyperparameter $t\geq 0$ that allows for weighting the two components separately. We use $t=2$, as $t>1$ gives tight lower bounds for RR. The adversary guesses on $c^{\prime}\%$ of the samples with the highest absolute scores. We sweep $c^{\prime}\in\{1,2,\dots,100\}$ and report the highest $\epsilon$} achieved at 95% confidence, averaged over 100 repetitions of the game (resamplings of counterfactual labels). In Appendix 12, we show how we obtain similar results for a fixed $c^{\prime}$.
@@ 6.3 CIFAR-10 experiments
For CIFAR-10, we treat the image classes as sensitive labels. We train standard convolutional neural networks with varying privacy budgets ${\epsilon}\in\{1,10,\infty\}$ (see model accuracy in Table 4, Appendix 11). We then audit Label DP using our observational game and report results in Table 1.
6.4 Criteo experiments
The Criteo dataset contains user click-through data with demographic information encoded as 13 numerical features and 26 categorical features. A binary label indicates whether the user clicked on the ad. The distribution of the labels is highly imbalanced, with only 3% of positives (clicks). The overall dataset contains over 4 billion click log data points over a period of 24 days. We followed the same setup in wu2022does where 1 million data points are selected for each day. We divide the data into 80% for training, 4% for validation, and 16% for testing. Model performance is evaluated using the log-loss metric on the test set.
We train gradient boosting decision trees with the CatBoost library (dorogush2018catboost) with varying privacy budgets ${\epsilon}\in\{0.1,1,2,4,8,\infty\}$ and evaluate label privacy using observational privacy auditing (Table 2). Table 5 (Appendix 11) shows model performance under different Label DP algorithms and values $\epsilon$}.
@@ 6.4 Criteo experiments
Table 2: Criteo. Auditing Label DP algorithms for different $\epsilon$} with $\delta=10^{-5}$. As in wu2022does, for LP-1ST (domain prior) at ${\epsilon}\in\{0.1,1,2\}$ the training process did not produce meaningful outcomes.
@@ 6.5 Comparison with Existing Methods
We compare our observational auditing approach against traditional canary-based methods to demonstrate the effectiveness and practicality of our framework. More specifically, we evaluate against the lightweight difficulty calibration MIA in watson2022difficulty, where membership scores are adjusted to the difficulty of correctly classifying the target sample. For each canary datapoint, we set the calibrated membership scores as the difference in the loss between the target model and the reference model $M^{\prime}$. Fig. 1 shows how our method achieves similar auditing results when compared to MIA on the CIFAR-10 and Criteo datasets.
@@
We leave to future work a comparison with more computationally intensive methods, such as zarifzadeh2024low, which require training multiple auxiliary (shadow) models to achieve state-of-the-art attack performances. However, we compare our approach to that of zarifzadeh2024low for the special case of one single reference model in Appendix 12.
Gap between theoretical and empirical epsilon. Fig. 1 shows a large gap between the theoretical and empirical bounds on $\epsilon$, especially for larger values of $\epsilon$. This gap is due to several factors that have been previously explored in the literature: (1) the attacks are black-box, i.e., they do not have access to intermediate outputs of the training procedure such as gradients (nasr2021adversary), (2) the theoretical $\epsilon$ is an upper bound that is not necessarily tight, (3) $f$-DP auditing is not always tight, especially at higher values of $\epsilon$} (mahloujifar2025auditing).
Figure 2: Auditing randomized response when the adversary guesses on 0.1% of samples. The counterfactual labels are generated either from the ground-truth distribution or sampled according to the output of a logistic regression model.
@@ 6.6 Synthetic data and auditing randomized response
We empirically demonstrate the tightness of our Label DP auditing algorithm for Randomized Response using synthetic data. The distribution consists of $k$ balanced classes, each generated from a $5$-dimensional Gaussian with the same covariance but a shifted mean. We generate $n=10^{6}$ samples and vary $k\in\{2,5,10\}$. Counterfactual labels $y^{1}$ are generated using either the true distribution or the predictions of a logistic regression model. Figure 2 shows empirical epsilon lower bounds at $95\%$ confidence when the adversary makes $0.1\%$ non-abstaining guesses. We obtain tight lower bounds for ${\epsilon}\in[1,4]$. At lower epsilons, the audit overestimates privacy loss due to a higher variance induced by a small number of guesses. Using more guesses at lower epsilon fixes the issue (Appendix 10).